The diagonal measure of a rectangular computer monitor screen is 53 cm. If the screen is 45 cm wide, then how tall is it?
Answer by oscargut(2103)

if t=tall using pythagoras
53^2=45^2+t^2
t^2=784
then
t=sqrt(784)=28
Answer: tall is 28 cm